A grandson tells his grandfather when I am at your age you will be 96 years old by then .the grandson is 8 years old now how old is his grandfather?
Answer by htmentor(1343)

Let G = the grandfather's current age
Let n = the number of years until the grandson is the grandfather's current age
We have two equations:
G + x = 96
8 + x = G
Subtract the 2nd from the 1st:
G - 8 = 96 - G
2G = 104
G = 52